OpenStudy (anonymous):

Solve for y: x=(y+3)/(2-y)

OpenStudy (anonymous):

x(2-y) = y + 3 2x - xy - y = 3 xy + y = 2x - 3 y(x + 1) = 2x - 3 y = (2x - 3)/(x + 1)

OpenStudy (anonymous):

Amazing. Thank you